# Seeing, Listening, Remembering, and Reasoning: A Multimodal Agent with Long-Term Memory

### v2

@@ 5.5 Case Study
Memorization Table LABEL:table:case_study_memory_webLABEL:table:case_study_memory_robot (§ 14) present two examples illustrating the episodic and semantic memories generated during memory access. Compared to memory-gemini-prompt, memory-7b-sft demonstrates (1) more detailed episodic memory generation, including richer scene descriptions, character actions and expressions, and dialogue; (2) improved recognition of identity equivalence, enabling consistent long-term tacking of human identities; and (3) richer semantic memory extraction, proactively generating knowledge about characters and environments.
Control To illustrate the control process in detail, Table LABEL:control_trajectory (§ 14) presents a complete generation trajectory of control-32b-rl. The input question is: "Is Tomasz a person with rich imagination or someone who lacks imagination?"
@@ 7 Acknowledgment
We would like to thank Xiran Suo, Wanjun Wang, and Liu Ding of ByteDance for their help with data annotation, and Peng Lin for creating the illustration.

### v4

@@ 5.5 Case Study
Memorization Table LABEL:table:case_study_memory_web, LABEL:table:case_study_memory_robot14) present two examples illustrating the episodic and semantic memories generated during memory access. Compared to memory-gemini-prompt, memory-7b-sft demonstrates (1) more detailed episodic memory generation, including richer scene descriptions, character actions and expressions, and dialogue; (2) improved recognition of identity equivalence, enabling consistent long-term tacking of human identities; and (3) richer semantic memory extraction, proactively generating knowledge about characters and environments.
Control To illustrate the control process in detail, Table LABEL:control_trajectory14) presents a complete generation trajectory of control-32b-rl. The input question is: "Is Tomasz a person with rich imagination or someone who lacks imagination?"
@@ 7 Acknowledgment
We would like to thank Xiran Suo, Wanjun Wang, Liu Ding, and Jianghui Xie of ByteDance for their help with data annotation, and Peng Lin for creating the illustration.